OpenStudy (nosnip):

solutions to 1/2 cosx=-0.5 (using cos graph)

OpenStudy (richyw):

Using the graph? here it is.

OpenStudy (richyw):

\[\frac{1}{2}\cos x = -0.5\]\[\cos{x}=-1\]so clearly\[x = (2n-1)\pi,\quad\pi\in\mathbb{Z}\]

OpenStudy (richyw):

which you can get by looking at the graph.

OpenStudy (richyw):

OOPS haha. obviously pi is not an integer. last line should read \[x = (2n-1)\pi,\quad n\in\mathbb{Z}\]

OpenStudy (richyw):

which just means \(n= 0,\pm 1,\pm 2, \pm3 \dots\)

OpenStudy (nosnip):

is there a way to do that without using pie?